OpenStudy (anonymous):

write as a single question 3+x/5 tell me im not as dumb as i feel

OpenStudy (lgbasallote):

you're not as dumb as you feel

jimthompson5910 (jim_thompson5910):

The question is \[\Large 3+\frac{x}{5}\] correct?

OpenStudy (anonymous):

hahaha nice thanks. not what i meant though xD

OpenStudy (anonymous):

and yeah jim thats the question. im supposed to write it as a single fraction

jimthompson5910 (jim_thompson5910):

alright, you can write 3 as 3/1

jimthompson5910 (jim_thompson5910):

What can you multiply that '1' by to get it to the LCD 5 ?

OpenStudy (anonymous):

5? right?

jimthompson5910 (jim_thompson5910):

yes, so do the same to the numerator as well to balance things out

OpenStudy (anonymous):

so itd be 15/5?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

so when i add the two fractions it should be 15x/5?

jimthompson5910 (jim_thompson5910):

so \[\Large 3+\frac{x}{5}\] becomes \[\Large \frac{15}{5}+\frac{x}{5}\]

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

From here, just add the numerators and place that over the common denominator

OpenStudy (anonymous):

so would it be 15+x/5?

jimthompson5910 (jim_thompson5910):

yes \[\Large \frac{15+x}{5}\]

jimthompson5910 (jim_thompson5910):

type (15+x)/5 to mean \[\Large \frac{15+x}{5}\]

OpenStudy (anonymous):

aaaaaaahhhhh yeah. i love this site :D

jimthompson5910 (jim_thompson5910):

here's the full step by step picture of what's happening \[\Large 3+\frac{x}{5}\] \[\Large \frac{3}{1}+\frac{x}{5}\] \[\Large \left(\frac{3}{1}\right)\left(\frac{5}{5} \right )+\frac{x}{5}\] \[\Large \frac{3 \times 5}{1\times 5}+\frac{x}{5}\] \[\Large \frac{15}{5}+\frac{x}{5}\] \[\Large \frac{15+x}{5}\] ----------------------------- So \[\Large 3+\frac{x}{5}\] simplifies to \[\Large \frac{15+x}{5}\]

OpenStudy (anonymous):

thanks :D

jimthompson5910 (jim_thompson5910):

That's just for reference

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

what if the variable is the denominator? like 1+3/x

jimthompson5910 (jim_thompson5910):

You do the same 1 = 1/1 then multiply both top and bottom by the LCD 'x'

jimthompson5910 (jim_thompson5910):

so \[\Large 1+\frac{3}{x}\] \[\Large \frac{1}{1}+\frac{3}{x}\] \[\Large \frac{x}{x}+\frac{3}{x}\]

OpenStudy (anonymous):

so x+3/x?

jimthompson5910 (jim_thompson5910):

type (x+3)/x to signify that all of "x+3" is over the 'x', but yes, it's \[\Large \frac{x+3}{x}\]

OpenStudy (anonymous):

oh ok. and thanks again

jimthompson5910 (jim_thompson5910):

keep in mind x+3/x means \[\Large x+\frac{3}{x}\] ------------------ while (x+3)/x means \[\Large \frac{x+3}{x}\]

OpenStudy (anonymous):

would 3+(x-2)/2 equate to (x+4)/2

jimthompson5910 (jim_thompson5910):

that is 100% correct

jimthompson5910 (jim_thompson5910):

nice work

OpenStudy (anonymous):

yes thank you. you explain it a lot better than my teachers xD

jimthompson5910 (jim_thompson5910):

I'm glad to hear that lol

jimthompson5910 (jim_thompson5910):

you're welcome